OpenStudy (anonymous):

Solve 6^x = 24 Solve 7.7^4x + 3 = 10.2

OpenStudy (saifoo.khan):

Solve 6^x = 24 \[\large x = \log_6 24\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x ≈ 30.84 x ≈ 1.77 x ≈ 0.56 x ≈ 0.23 choices

OpenStudy (saifoo.khan):

Do you have a calc?

OpenStudy (anonymous):

Yeah

OpenStudy (saifoo.khan):

insert the values like this: \[\Large \log_6 24 = \frac{\log 24}{\log 6}\]

OpenStudy (anonymous):

Can you show me

OpenStudy (saifoo.khan):

Sure. Which calc do you have?

OpenStudy (anonymous):

basic

OpenStudy (saifoo.khan):

Do you know it's model # ?

OpenStudy (anonymous):

no why

OpenStudy (saifoo.khan):

If it's like the first one, then it wont help here.. If it's like the second one, then it will help.

OpenStudy (anonymous):

well can you just show me how to do it

OpenStudy (saifoo.khan):

You may use this: http://www.wolframalpha.com/input/?i=log6+24&dataset=